OpenStudy (anonymous):

4x-2y=7 in slope intercept form. I keep getting a fraction for y

OpenStudy (anonymous):

could you show me your work? what you got?

OpenStudy (anonymous):

4x-2y=7 -2y=-4x+7 -2y=-4/2x+7/2 y=-4/2x+7/2

OpenStudy (anonymous):

Y=mx+b

OpenStudy (anonymous):

you almost have it. -2y = -4x + 7 Now divide both sides by -2. This leaves y by itself. -4 divided by -2 is ___. You would simplify leave 7 divided by -2 alone. The y-intercept (b) is -(7/2)

OpenStudy (anonymous):

ok so the slope would be 2/1 and b= 7/2 so i would just plot that as a point or i have to find 7/2 on my graph

OpenStudy (anonymous):

as a fraction

OpenStudy (anonymous):

yes in slope-intercept form it would be: y = 2x - (7/2). Do you actually have to graph it or just find it in slope-intercept form?

OpenStudy (anonymous):

no i dont have to graph i was given two lines and i have to decide it they are parallel perpendicular, or neither the other line is y=-4x+2

OpenStudy (anonymous):

its neither right?

OpenStudy (anonymous):

tell me why you think its neither

OpenStudy (anonymous):

so if they are parallel they have the same slope? i cant remember what the parameters were for them being perpendicular

OpenStudy (anonymous):

Im a little rusty

OpenStudy (anonymous):

ok. do you have to draw it by hand or are you allowed to graph it on a calculator?

OpenStudy (anonymous):

I dont have to draw it i just have to state what it is

OpenStudy (anonymous):

Ok, by graphing it on a calculator, the lines intersect meaning its perpendicular.